OpenStudy (babyslapmafro):

Please help me find the Laplace transform of the given expression that has the properties y(0)=1 and y'(0)=-1: y"+3y'

OpenStudy (anonymous):

is the eq.y"+3y'=0

OpenStudy (babyslapmafro):

It is not an equation, just an expression

OpenStudy (ybarrap):

$$ \mathcal{L}\left\{y'(t)\right\}(s)=s Y(s) - y(0) \\ \mathcal{L}\left\{y''(t)\right\}(s)= s^2 Y(s) - s y(0) - y'(0) \\ $$ So, taking Laplace of y''+3y': $$ \mathcal{L}\left\{y"\!+~3y'\right\}(s)= s^2 Y(s) - s y(0) - y'(0) +3\left(s Y(s) - y(0) \right )\\ =Y(s)(s^2+3s)-s+1-3s\\ =Y(s)(s^2+3s)-4s+1$$

OpenStudy (babyslapmafro):

Thanks for taking the time to do that but the answer in the book is: \[(s^2+3s)F(s)-s-2\]

OpenStudy (ybarrap):

see typo - $$ \mathcal{L}\left\{y"\!+~3y'\right\}(s)= s^2 Y(s) - s y(0) - y'(0) +3\left(s Y(s) - y(0) \right )\\ =Y(s)(s^2+3s)-s+1-\color{red}{3}\\ =Y(s)(s^2+3s)-s-2 $$

OpenStudy (babyslapmafro):

oh ok i see, thanks for working that out...

OpenStudy (ybarrap):

np